OpenStudy (zenmo):

(Sketching polar graphs) sketch the curve with the given polar equation by first sketching the graph of r as a function of theta in Cartesian coordinates.

OpenStudy (zenmo):

\[r = 2 + 2\cos \theta \]

OpenStudy (zenmo):

I am not sure how to start this, but I got the general idea of the problem.

OpenStudy (phi):

pick some values of theta, and generate a table for theta, r values. plot the points connect the dots.

OpenStudy (zenmo):

what theta should I use?

OpenStudy (phi):

cos theta has a period of 2pi (360 deg) so we only need angles between 0 and 2pi I would use 0,30,60,90 as a first step.

OpenStudy (zenmo):

Yea choosing pi/6, pi/4, pi/3, pi/2, 2pi/3 , 3pi/4, 5pi/6, pi, so I would need to continue from pi to 2pi to get the picture?